# Bounds on the realizations of zero-nonzero patterns and sign conditions of polynomials restricted to varieties and applications

### v2

@@ 3.2. Bound on the ${\varepsilon}$-entropy of real varieties
A quantitative result from real (or more accurately metric) algebraic geometry that is now of great interest in applications (in problems originating in machine learning) is a tight upper bound on the volume of an ${\varepsilon}$-tube around a real algebraic variety $V$ intersected with a ball of a given radius. A recent result in this direction is [9, Theorem 1.1] which gives a tight upper bound on this volume depending on the degree $d$ of the polynomials defining $V$, the codimension of $V$.
@@
The following result appears in [51, page 4, Table 1, entry on real varieties].
Theorem H.
@@
Note that right-hand side of the inequality in Theorem 10 depends very weakly on the embedding dimension $N$ ($\log N$ instead of $N$) compared to the bound in Theorem H. Notice also that while the bound in Theorem H is in terms of the degrees of the real polynomials cutting out $V$, the bound in Theorem 10 is in terms of the degree of $V$ (as a complex variety) which is a more intrinsic invariant. It is well known that any affine variety of degree $D$ in $\mathrm{C}^{N}$ can be cut out by at most $N+1$ polynomials of degree at most $D$. However, if we use this fact in conjunction with Theorem H, we will end up with a much weaker bound than the one in Theorem 10 ($N$ instead of $\log N$ in the expression on the right-hand side).
@@ 3.3. Lower bounds for algebraic computation trees
Note that the above lower bounds are independent of the ambient dimension $N$. As in previous remarks this is useful in situations where $\dim V$ and $\log\deg V$ are very small compared to the dimension of the ambient space.
@@
For example, consider the Grassmannian variety $\mathrm{Gr}(m,N)$ of $m$-dimensional subspaces in $\mathrm{C}^{N}$, as a subvariety of $\mathbb{P}_{\mathrm{C}}^{\binom{N}{m}-1}$. Let $V=V_{m,N}\subset\mathrm{C}^{\binom{N}{m}}$ be the affine cone over $\mathrm{Gr}(m,N)$. Then, it is well known (see for example [33, Corollary 3.2.14]) that
@@ 3.4. Relative ranks in vector spaces and algebras
The notion of rank of tensors play a very important role in many applications – from understanding the complexity of matrix multiplication, to machine learning and quantum information theory (see for example [30] for an overview). We introduce here a notion of relative rank for finite dimensional vector spaces and algebras (relative to an arbitrary algebraic subset $\Delta$ of the vector space or algebra) which generalizes the notion of rank for tensors. We prove lower bounds on the relative rank for algebraically defined subsets $W$ of the vector space or algebra, in terms of the dimension and degree of $\Delta$ and the cardinality (or the number of connected components) of the set (Theorems 12, 13, 14 and 15 below).
Note that while proving upper bounds on the classical rank of specific tensors (such as the matrix multiplication tensor) [30] is an important and active area of research, proving lower bounds on the rank of sets of tensors (particularly in the relative setting) has not been equally well studied. We believe the results stated below will prove useful in applications in addition to quantum complexity theory that we discuss here.
@@ 3.4.1. Relative rank in vector spaces
Let $k$ be an algebraically closed field and $V$ a finite dimensional vector space. Let $\Delta\subset V$ be an algebraic subset with $\dim(\Delta)=p$ and $\deg(\Delta)=D$, and $\mathcal{P}\subset k[V]_{\leq\delta}$ with $\mathrm{card}(\mathcal{P})=s$. Then, for any finite $\mathcal{P}$-constructible subset $W\subset A$
@@
In the real closed case we have a slightly worse lower bound (with an added non-singularity assumption on $V$).
@@
$\mathrm{rank}_{V,\Delta}(W)\in\Omega\left(\frac{\log(\mathrm{card}(W))}{p(\log\delta+\log s+\log D)}\right)$
@@ 3.4.2. Relative rank in algebras
$\mathrm{mmrank}_{A,\Delta}(W)\in\Omega\left(\frac{\log(b_{0}(W))}{p(\log\delta+\log s+\log\log b_{0}(W)+\log D)}\right)$
@@
3.4.3. Example application of Theorem 12
Example 3.4 (Ranks of tensors).
Let $k$ be algebraically closed and let $V_{1},\ldots,V_{m}$ be finite dimensional vector spaces with $\dim V_{i}=N_{i}$. Let $V=V_{1}\otimes\cdots\otimes V_{m}$. For $1\leq i\leq m$, let $B_{i}=(e_{i,j})_{1\leq j\leq N_{i}}$ be a basis of $V_{i}$, and $B=B_{1}\otimes\cdots\otimes B_{m}$ the induced basis of $V$. The coordinates of an element $v$ with respect to the basis $B$ is then given by a $m$-dimensional array in $k^{N_{1}\times\cdots\times N_{m}}$. Let $B^{\prime}\subset B$ be a fixed set and $W\subset V$, be defined as the finite subset of elements of $V$, whose coordinates with respect to the basis elements in $B^{\prime}$ belong to $\{0,1\}$ and whose coordinates with respect to the basis elements $B-B^{\prime}$ are 00. Clearly, $\mathrm{card}(W)=2^{\mathrm{card}(B^{\prime})}$. (In terms of arrays $W$ is a finite set of $N_{1}\times\cdots\times N_{m}$ arrays whose entries in the positions indexed by $B^{\prime}$ are 00 or $1$ and whose remaining entries are 00).
Let $\Delta\subset V$ be the set of tensors of rank $\leq 1$. Then, $\Delta$ can be identified with the affine cone over the image of the Segre embedding
$\mathbb{P}_{k}^{N_{1}-1}\times\cdots\mathbb{P}_{k}^{N_{m}-1}\rightarrow\mathbb{P}_{k}^{N_{1}\cdots N_{m}-1},$
which yields:
$p=\dim\Delta=N_{1}+\cdots+N_{m}-m+1,$
and
$D=\deg\Delta=\binom{N_{1}+\cdots+N_{m}-m}{N_{1}-1,\ldots,N_{m}-1}$
using the multinomial notation (see for example [16, Proposition 2.11]). Applying Theorem 12, and noting that $\mathrm{card}(W)=2^{\mathrm{card}(B^{\prime})}$, we obtain
$\mathrm{rank}_{V,\Delta}(W)=\Omega\left(\frac{\mathrm{card}(B^{\prime})}{p\cdot(\log(N_{1}\cdots N_{m})+\log D)}\right).$
Taking $m=2$, and $N_{1}=N_{2}=N$, and $B^{\prime}\subset B$ with $\mathrm{card}(B^{\prime})=N^{1+{\varepsilon}}$, for some ${\varepsilon},0<{\varepsilon}<1$, we obtain that
$\mathrm{rank}_{V,\Delta}(W)=\Omega\left(\frac{N^{{\varepsilon}}}{\log N}\right).$
Thus, in any set of $N\times N$ zero-nonzero matrices with non-zero entries allowed only in a fixed set of $N^{1+{\varepsilon}}$ positions, there exists a matrix of rank at least $C\cdot\frac{N^{\varepsilon}}{\log N}$, where $C>0$ is an absolute constant.
Note that a possible alternative approach to proving the above statement is to use the well known Kővári-Sós-Turán (KST) theorem (see for example [52, Theorem 1.4.2]). Identifying the set $B$ with $[1,N]\times[1,N]$, if the chosen set of positions, $B^{\prime}\subset B$ contains a “combinatorial square” of size $s\times s$ (i.e. $I\times J$, with $\mathrm{card}(I)=\mathrm{card}(J)=s,I,J\subset[1,N]$), then clearly, $\mathrm{rank}_{V,\Delta}(W)\geq s$. The KST theorem guarantees the existence of such a square if $\mathrm{card}(B^{\prime})\geq C(s)\cdot N^{2-1/s}$. In the KST theorem $s$ is constant. For $s$ growing with $N$, the cardinality of $B^{\prime}$ needs to be even larger. Thus, for $s=N^{\varepsilon}/\log N$, $\mathrm{card}(B^{\prime})$ needs to be at least $N^{2-o(1)}$ to be guaranteed that $B^{\prime}$ contains a combinatorial square with $s=N^{\varepsilon}$. However, $N^{1+{\varepsilon}}\ll N^{2-o(1)}$, and thus this approach cannot be used to prove the claim above.
3.4.4. Example of application of Theorem 14
Example 3.5 (Lengths of permutations).
Let $V$ be an $n$-dimensional $k$-vector space and $A=\mathrm{End}(V)$. Let $B=\{v_{1},\ldots,v_{n}\}$ be a vector space basis of $A$, and $W$ the subset of $A$ of those endomorphism that permutes the elements of $B$. Then, $W$ is a zero-dimensional algebraic subset of $A$ cut out by $n^{2}+2n$ equations of degrees at most $2$ ($n^{2}$ equations of degree $2$ and $2n$ equations of degree $1$), and $\mathrm{card}(W)=n!$.
If $\Delta\subset A$ be any $p$-dimensional variety having degree $D$, then using Theorem 1 we have:
$\mathrm{rank}_{A,\Delta}(W)=\Omega\left(\frac{n\log n}{(p+1)\log n+\log D}\right).$
This bound is asymptotically tight.
Consider for example $A=\mathrm{End}(V)$, with $V$ a $k$-vector space with basis $v_{1},\ldots,v_{n}$, and $\Delta$ the union of $\mathrm{Id}_{V}$, and the endomorphisms $\pi_{i},1\leq i\leq n-1$, where $\pi_{i}$ acts on $B$ by transposing $v_{i},v_{i+1}$ and by the identity on $v_{j},j\not\in\{i,i+1\}$. Then as is well known each element of $W$ can be expressed as a product of at most $2N$ elements of $\mathrm{Id}_{V},\pi_{1},\ldots,\pi_{N-1}$.
So, $\mathrm{rank}_{A,\Delta}(W)=O(n)$, while in the lower bound above we have $p=0$, $D=n$, which gives $\mathrm{rank}_{A,\Delta}(W)=\Omega(n)$. Indeed the $\Omega(n)$ lower bound is valid for $p=O(1)$ and $D\leq n$.
@@ 3.5.1. Shannon’s lower bound for classical circuits
Boolean circuit complexity measures complexity of Boolean functions $f:\{0,1\}^{n}\rightarrow\{0,1\}$ by the size of the smallest circuit computing $f$. We assume that the gates of a circuit come from some fixed finite universal set of gates. We refer the reader to [49, Page 7, Definition 3.1] for the precise definition of a circuit. Notice that in this case the set of circuits of bounded size is finite.
Shannon [40] showed using a counting argument (using the fact that as noted before that the set of circuits of bounded size is finite) very early that almost all Boolean functions need circuits of size $\Omega({2^{n}}/{n})$ (see [49, Page 90, Theorem 2.1]. Here almost all refers to the fact that the number of Boolean functions that require circuits of size $\Omega({2^{n}}/{n})$ is bounded from below by $2^{2^{n}}(1-o(1))$. It is also known (see for instance [49, Page 92, Theorem 2.2]) that every Boolean function $f:\{0,1\}^{n}\rightarrow\{0,1\}$ can be computed by a circuit of size $O({2^{n}}/{n})$.
@@ 3.5.2. Analog of Shannon’s lower bound for quantum circuits
A $q$-ary quantum gate $g$ is an element $U_{g}\in\mathbf{U}(V^{\otimes q})$. In a quantum circuit $C$, with input $n$-qubits, such a gate operates on a subset of $q$ of the $n$-qubits (say labelled by the tuple $\mathbf{i}=(i_{1},\ldots,i_{q}),1\leq i_{1}<\cdots<i_{q}\leq$), which gives rise to an element $U_{g,\mathbf{i}}$ of $\mathbf{U}(V^{\otimes n})$ which can be identified with $U_{g}\otimes\mathrm{Id}_{V^{\otimes(n-q)}}$. (after rearranging the order of the copies of $V$ in the tensor power $V^{\otimes n}$ bringing the ones indexed by $(i_{1},\ldots,i_{q})$ to the the positions $(1,\ldots,q)$).
@@
$\Delta=\Delta_{n,q}=\{U_{g,\mathbf{i}}\mid U_{g}\in\mathbf{U}(V^{\otimes q}),\mathbf{i}=(i_{1},\ldots,i_{q}),1\leq i_{1}<\cdots<i_{q}\},$
@@
For arbitrary $\Delta\subset\mathrm{End}(V^{\otimes n})$ we will call $\mathrm{rank}_{\mathrm{End}(V^{\otimes n}),\Delta}(U)$ the quantum complexity of $U$ relative to $\Delta$.
@@
In analogy with classical complexity, $\Delta$ could be thought of as a set of quantum oracles – with unit cost for each application of an element of $\Delta$ in an quantum circuit. Notice though we do not assume that $\Delta$ is contained in the unitary group $\mathbf{U}(V^{\otimes n})$, but it is allowed to be.
@@
In the following theorem we consider $A=\mathrm{End}(V^{\otimes(n+t+1)})$ as a real subspace of the $4^{n+t+1}\times 4^{n+t+1}$ dimensional real vector space $A_{\mathbb{R}}=\mathrm{End}(V_{\mathbb{R}}^{\otimes(n+1+1)})$ considering $V$ as a $4$-dimensional $\mathbb{R}$-vector space denoted $V_{\mathbb{R}}$.
@@
$C\cdot\frac{2^{n}}{p\cdot(n+t+\log D)}.$
@@
[37, Theorem 7.1] For each $q\geq 1$, there exists $C=C_{q},0<c<1$ (depending only on $q$), such that for each $n>0$ there exists Boolean functions $f:\{0,1\}^{n}\rightarrow\{0,1\}$, such that the quantum circuit complexity of $f$ with $t+1$ ancillary bits is at least $C\cdot\frac{2^{n}}{n+t}$.
@@
This lower bound goes to 00 as $\delta\rightarrow 0$. So it does not produce any meaningful lower bound for $\delta=0$ (which is the case in Definition 3.10). We also note that there is recent work in constructing explicit Boolean functions with exponential quantum circuit complexity using sophisticated tools from transcendental number theory [26]. Our rank argument does not produce such functions.

### v3

@@ 3.1. Bound on the ${\varepsilon}$-entropy of real varieties
A quantitative result from real (or more accurately metric) algebraic geometry that is now of great interest in applications (in problems originating in machine learning) is a tight upper bound on the volume of an ${\varepsilon}$-tube around a real algebraic variety $V$ intersected with a ball of a given radius. A recent result in this direction is [7, Theorem 1.1] which gives a tight upper bound on this volume depending on the degree $d$ of the polynomials defining $V$ and the codimension of $V$.
@@
The following result appears in [47, page 4, Table 1, entry on real varieties].
Theorem G.
@@
Remark 13.
Note that right-hand side of the inequality in Theorem 8 depends very weakly on the embedding dimension $N$ ($\log N$ instead of $N$) compared to the bound in Theorem G. Notice also that while the bound in Theorem G is in terms of the degrees of the real polynomials cutting out $V$, the bound in Theorem 8 is in terms of the degree of $V$ (as a complex variety) which is a more intrinsic invariant. It is well known that any affine variety of degree $D$ in $\mathrm{C}^{N}$ can be cut out by at most $N+1$ polynomials of degree at most $D$. However, if we use this fact in conjunction with Theorem G, we will end up with a much weaker bound than the one in Theorem 8 ($N$ instead of $\log N$ in the expression on the right-hand side).
Remark 14.
Using Remark 9 the right hand side of both inequalities in Theorem 8 can be replaced by
$p\cdot\log(1/{\varepsilon})+O(p\cdot(\log\deg(V)+\log(\deg(V)+p)),$
which is completely independent of $N$.
@@ 3.2. Lower bounds for algebraic computation trees
Note that the above lower bounds are independent of the ambient dimension $N$. This is useful in situations where $\dim V$ and $\log\deg V$ are very small compared to the dimension of the ambient space.
@@
For example, consider the Grassmannian variety $\mathrm{Gr}(m,N)$ of $m$-dimensional subspaces in $\mathrm{C}^{N}$, as a subvariety of $\mathbb{P}_{\mathrm{C}}^{\binom{N}{m}-1}$. Let $V=V_{m,N}\subset\mathrm{C}^{\binom{N}{m}}$ be the affine cone over $\mathrm{Gr}(m,N)$. Then, it is well known (see for example [30, Corollary 3.2.14]) that
@@ 3.3. Relative ranks in vector spaces and algebras
The notion of rank of tensors play a very important role in many applications – from understanding the complexity of matrix multiplication, to machine learning and quantum information theory (see for example [27] for an overview). We introduce here a notion of relative rank for finite dimensional vector spaces and algebras (relative to an arbitrary algebraic subset $\Delta$ of the vector space or algebra) which generalizes the notion of rank for tensors. We prove lower bounds on the relative rank for algebraically defined subsets $W$ of the vector space or algebra, in terms of the dimension and degree of $\Delta$ and the cardinality (or the number of connected components) of the set (Theorems 10, 11, 12 and 13 below).
Note that while proving upper bounds on the classical rank of specific tensors, such as the matrix multiplication tensor (see for example [26, 27] is an important and active area of research, proving lower bounds on the rank of sets of tensors (particularly in the relative setting) has not been equally well studied. We believe the results stated below will prove useful in applications in addition to quantum complexity theory that we discuss here.
@@ 3.3.1. Relative rank in vector spaces
Let $k$ be an algebraically closed field and $V$ a finite dimensional vector space. Let $\Delta\subset V$ be an algebraic subset with $\dim(\Delta)=p$ and $\deg(\Delta)=D$, and $\mathcal{P}\subset k[V]_{\leq\delta}$ with $\mathrm{card}(\mathcal{P})=s$. Then, for any finite $\mathcal{P}$-constructible subset $W\subset V$
@@
In the real closed case we have a similar lower bound (with an added non-singularity assumption on $V$).
@@ 3.3.2. Relative rank in algebras
$\mathrm{mmrank}_{A,\Delta}(W)\in\Omega\left(\frac{\log(b_{0}(W))}{p(\log\delta+\log s+\log\log b_{0}(W))+\log D}\right)$
@@ 3.4.1. Shannon’s lower bound for classical circuits
Boolean circuit complexity measures complexity of Boolean functions $f:\{0,1\}^{n}\rightarrow\{0,1\}$ by the size of the smallest circuit computing $f$. We assume that the gates of a circuit come from some fixed finite universal set of gates. We refer the reader to [45, Page 7, Definition 3.1] for the precise definition of a circuit. Notice that in this case the set of circuits of bounded size is finite.
Shannon [37] showed using a counting argument (using the fact that as noted before that the set of circuits of bounded size is finite) very early that almost all Boolean functions need circuits of size $\Omega({2^{n}}/{n})$ (see [45, Page 90, Theorem 2.1]. Here almost all refers to the fact that the number of Boolean functions that require circuits of size $\Omega({2^{n}}/{n})$ is bounded from below by $2^{2^{n}}(1-o(1))$. It is also known (see for instance [45, Page 92, Theorem 2.2]) that every Boolean function $f:\{0,1\}^{n}\rightarrow\{0,1\}$ can be computed by a circuit of size $O({2^{n}}/{n})$.
@@ 3.4.2. Analog of Shannon’s lower bound for quantum circuits
A $q$-ary quantum gate $g$ is an element $U_{g}\in\mathbf{U}(V^{\otimes q})$. In a quantum circuit $C$, with input $n$-qubits, such a gate operates on a subset of $q$ of the $n$-qubits (say labelled by the tuple $\mathbf{i}=(i_{1},\ldots,i_{q}),1\leq i_{1}<\cdots<i_{q}\leq N$), which gives rise to an element $U_{g,\mathbf{i}}$ of $\mathbf{U}(V^{\otimes n})$ which can be identified with $U_{g}\otimes\mathrm{Id}_{V^{\otimes(n-q)}}$. (after rearranging the order of the copies of $V$ in the tensor power $V^{\otimes n}$ bringing the ones indexed by $(i_{1},\ldots,i_{q})$ to the the positions $(1,\ldots,q)$).
@@
$\Delta=\Delta_{n,q}=\{U_{g,\mathbf{i}}\mid U_{g}\in\mathbf{U}(V^{\otimes q}),\mathbf{i}=(i_{1},\ldots,i_{q}),1\leq i_{1}<\cdots<i_{q}\leq n\},$
@@
For arbitrary $\Delta\subset\mathrm{End}(V^{\otimes n})$ we will call
$\mathrm{rank}_{\mathrm{End}(V^{\otimes n}),\Delta}(U)$
the quantum complexity of $U$ relative to $\Delta$.
@@
In analogy with classical complexity, $\Delta$ could be thought of as an (quantum) oracle – with unit cost for each application of an element of $\Delta$ in an quantum circuit. Notice though we do not assume that $\Delta$ is contained in the unitary group $\mathbf{U}(V^{\otimes n})$.
Remark 22 (Quantum complexity classes with oracles).
We note here that complexity results relative to oracles play an important role in classical as well as quantum complexity. For example, it has been shown that relative to an oracle $\mathbf{BQP}$ is separated from the classical polynomial hierarchy $\mathbf{PH}$ [32]. In another direction Aaronson [1] showed that by extending the set of gates from unitary to arbitrary linear maps the corresponding “quantum” complexity class that one obtains which is analogous to the well-known quantum complexity class $\mathbf{BQP}$ is equal to the classical complexity class Probabilistic Polynomial Time $\mathbf{PP}$. Extending the set of gates corresponds in the notation introduced in this paper (Definition 3.5) letting
$\Delta^{\prime}=\Delta^{\prime}_{n,q}=\{L_{g,\mathbf{i}}\mid L_{g}\in\mathrm{End}(V^{\otimes q}),\mathbf{i}=(i_{1},\ldots,i_{q}),1\leq i_{1}<\cdots<i_{q}\leq n\},$
and defining the complexity of $U\in\mathbf{U}(V^{\otimes n})$ to be
$\mathrm{rank}_{\mathrm{End}(V^{\otimes n}),\Delta^{\prime}}(U).$
Thus, clearly a small increase in the set $\Delta$ in Definition 3.5 can lead to a substantial increase in the computational power of quantum circuits (assuming the widely believed hypothesis that $\mathbf{BQP}$ is properly contained in $\mathbf{PP}$).
The main theorems in this section (Theorems 14 and 15) can be interpreted as pointing in the opposite direction (from the point of view Shannon’s lower bound theorem). They imply that Shannon’s lower bound theorem holds true for quantum circuits even relative to oracles much larger than $\Delta$ in Definition 3.5. We can allow $\Delta$ to be arbitrary unions of non-singular real complete intersections, of dimension bounded by some constant (independent of $n$), and degree bounded by $O(1)^{n}$ (exponential in $n$). Such a $\Delta$ can also allow for quantum gates that are non-local (i.e.acting on a non-constant number of qubits). In comparison, $\Delta_{n,q}$ in Definition 3.5, is contained in a union of real complete intersections of dimension $\leq q^{2}=O_{q}(1)$ and degree $\sum_{i=1}^{q}\binom{n}{q}=O(n)^{q}$ (polynomial in $n$) and each gate acts on at most $q$ qubits.
@@
In the following theorem we consider $A=\mathrm{End}(V^{\otimes(n+t+1)})$ as a real subspace of the $4^{n+t+1}\times 4^{n+t+1}$ dimensional real vector space $A_{\mathbb{R}}=\mathrm{End}(V_{\mathbb{R}}^{\otimes(n+t+1)})$ considering $V$ as a $4$-dimensional $\mathbb{R}$-vector space denoted $V_{\mathbb{R}}$.
@@
$C\cdot\frac{2^{n}}{p\cdot(n+t)+\log D}.$
@@
[34, Theorem 7.1] For each $q\geq 1$, there exists $C=C_{q},0<c<1$ (depending only on $q$), such that for each $n>0$ there exists Boolean functions $f:\{0,1\}^{n}\rightarrow\{0,1\}$, such that the quantum circuit complexity of $f$ with $t+1$ ancillary bits is at least $C\cdot\frac{2^{n}}{n+t}$.
@@
This lower bound goes to 00 as $\delta\rightarrow 0$. So it does not produce any meaningful lower bound for $\delta=0$ (which is the case in Definition 3.8). Note that having a positive $\delta$ or not in the definition of acceptance for a probabilistic complexity class (cf. Eqn. (3.4)) is a very important distinction. In classical complexity theory this distinction differentiates the complexity class $\mathbf{BPP}$ (with $\delta>0$) and the class $\mathbf{PP}$ (where $\delta=0$). As is well known,
$\mathbf{BPP}\subset\mathbf{NP}\subset\mathbf{PP},$
and both inclusions are believed to be proper.
There is recent work in constructing explicit Boolean functions with exponential quantum circuit complexity using sophisticated tools from transcendental number theory [22]. Our rank argument does not produce such functions.